# RE D shift: quasinormal modes of black holes embedded in halos of matter

### v1

@@ I Introduction
For decades, black holes (BHs) have only been considered as mathematical solutions to the Einstein field equations, without any astrophysical relevance. It took many years before they were accepted as anything other than a mathematical curiosity. Even though “there should be a law of nature to prevent a star from behaving in this absurd way” [^1] Arthur Eddington, Royal Astronomical Society conference, January 1935, we now are certain that BHs not only exist but also directly affect the way we understand the cosmos in its most extreme regimes. Recent experiments detected gravitational waves (GWs) from binary BHs and neutron stars [1, 2]. These detections are now characterized as the pinnacle of GW astronomy [3, 4, 5]. Combined with the foundational results of Gravity Probe A, B [6, 7], astrometry measurements at the center of our galaxy [8, 9], X-ray spectroscopy [10, 11, 12], infrared spectroscopy [13, 14] and radio astronomy [15, 16], the direct detection of GWs established the field of multimessenger astrophysics [17]. BHs do not only exist in nature, but also interact with other celestial objects, surrounding matter and the interstellar medium that veils the universe.
@@
Decades of simulations and observations have shown that any event involving a binary BH merger is likely to end in the same way; the GW amplitude of the merger will decay as a superposition of characteristic damped sinusoids. The damped oscillations are described by quasinormal modes (QNMs) [18, 19, 20], i.e. a complex set of frequencies which are an intrinsic property of dissipative systems. Calculating QNMs is achieved with the use of BH perturbation theory, and has led to the establishment of the BH spectroscopy program [21, 22, 23, 24, 25]. QNM frequencies are eigenmodes that are linked to the generic response of a BH under small fluctuations to the background spacetime and can help us to understand the nature of the remnant, providing precise measurements of its mass and spin, as well as allowing for tests of General Relativity (GR) [21, 19, 26, 27].
@@
The striking majority of QNM literature analyzes the spectral content of vacuum BHs, clearly an idealization. Massive BHs sit at the center of galaxies which are orders of magnitude more massive, where gaseous accretion disks [28, 29, 30], dense dark matter spikes [31, 32], and galactic dark matter halos form that affect the whole galaxy [33]. When two galaxies merge, baryonic and dark matter will interact and the supermassive BHs in their cores will meld into a new supermassive BH [34]; this event is expected to emit mHz GWs and falls into the frequency band of spaceborne interferometers, like the Laser Interferometer Space Antenna (LISA) [35, 36, 37, 38]. In this extreme case, the entire galaxy is an active participant in the dynamics and relaxation of the BH(s). The universe is also permeated with interstellar plasma, which prevents propagation of low-frequency electromagnetic waves, a screening mechanism occurs that entirely cloaks electromagnetic from GWs, which can have strong implications in BH spectroscopy [39].
@@
Taking into account astrophysical environments around BHs leads to interesting effects that can affect the emission and propagation of GWs. Recently, an interesting phenomenon – spectral instability – regained a lot of interest and is directly tied to the existence of astrophysical environments around BHs. Spectral instabilities occur when QNM excessively migrate in the complex plane. This happens when environmental fluctuations, such as thin matter shells or dark matter spikes, are introduced in the vicinity of vacuum BHs. It first appeared in Refs. [40, 41], and eventually led to the emergence of an extensive literature in relation to spectral instabilities [42, 43, 44, 45, 46, 47] and the use of non-modal tools to complement BH spectroscopy and QNMs, such as the pseudospectrum [48, 49, 50, 51, 52, 24, 53, 54, 55, 56, 57, 58, 59, 60]. Thus, treating compact objects in non-vacuum systems will provide us with information regarding astrophysical environments and their phenomenology in GW astronomy [61].
@@
In this work, we aim in giving an extended analysis of axial QNMs for numerically-constructed BH spacetimes with different matter profiles, such as the Hernquist [62] and Navarro-Frenk-White (NFW) distribution [63] that are suited for different kinds of environments. Our expectation is first to provide a public code that can treat axial QNM calculations when any form of matter distribution is present, and to inspect if the redshift effect found for BHs surrounded by Hernquist-type matter [25], when the compactness is varied, occurs for other matter profiles, thus extending the results regarding the detectability of astrophysical environments with axial QNMs [64]. In what follows, we use geometrized units such that $G=c=1$.
@@ II BHs in generic matter halos
Surrounding matter around perturbed astrophysical BHs can affect their GW emission and propagation. The research on non-vacuum BH QNMs [65, 66] is still in full bloom, with particular focus on galactic supermassive BHs [67, 68]. Most recently, an exact general-relativistic solution of the field equations was found that describes a Schwarzschild BH embedded in a matter halo, that poses the most massive part of any galaxy. Although this approach sounds simplistic, since it does not account for luminous matter and accretion disks, to date only one exact solution of such a static spacetime has been found, where the matter halo is described by the energy density of the Hernquist profile [62, 67]. In general, more spacetimes of the same nature have been constructed but only numerically or with various approximate techniques [69, 70, 71, 72, 73, 74, 75, 76, 77, 78]. These fully general-relativistic configurations can, overall, provide a good description of the geometry of galaxies harboring supermassive BHs and have been proven very useful to constrain the environmental surrounding of BHs [67]. In what follows we describe them briefly.
@@ II.1 Hernquist-type matter around static BHs
An exact, fully-relativistic solution that describes a BH immersed in a halo of matter was obtained recently [67]. The environmental contribution in this case is accounted for through the Einstein cluster approach, which describes the matter content in terms of an anisotropic stress-energy tensor, with only tangential ($P_{t}$) and vanishing radial pressure ($P_{r}=0$), i.e. $T^{\mu}{{}_{\nu}}=\text{diag}(-\rho,0,P_{t},P_{t})$. Choosing the energy density of the fluid to follow the Hernquist distribution [62], i.e.,
@@
where $M_{\textrm{BH}}$ is the BH mass. The factor $1-2M_{\rm BH}/r$ in the equations above is chosen to truncate the matter at the BH horizon. It can be chosen freely, in Eq. (1) at the expense of not having a closed-form solution for the spacetime geometry. Other truncation factors, such as $1-4M_{\rm BH}/r$ will mimic matter distributions which conform better with relativistic evolutions of dark matter profiles [79, 80].
@@
The BH spacetime has an event horizon at $r=r_{h}=2M_{\textrm{BH}}$ and a curvature singularity at $r=0$. Assuming $\xi>0$ and a hierarchy of scales, i.e. $M_{\textrm{BH}}\ll M\ll a_{0}$, ensures that there are no curvature singularities outside the horizon.
@@
with $U$ being the Newtonian potential of the matter distribution [81]. For the Hernquist distribution discussed here, $\Upsilon=2U$. We will use Eq. (6) to calculate the corrections to the light ring and its frequency, and determine the QNM frequencies in the eikonal regime.
@@ II.2 Static BHs surrounded by generic matter profiles: a numerical study
where $\rho_{0}$ is the energy density at $r=a_{0}$. The exponents $\beta$ and $\gamma$ control the slope of the distribution at large and small distances, respectively, while $\alpha$ determines the sharpness of the slope transition [82]. Equation (7) motivated Refs. [72, 75] to develop a numerical pipeline to constructs static, and spherically symmetric BH solutions embedded in spherically symmetric, but otherwise generic, astrophysical environments.
In this work, we employ the numerical approach developed in [72, 75] and analyze two specific combinations of the $(\alpha,\beta,\gamma)$ parameters from Eq. (7) in order to build numerical BH solutions: (i) $(\alpha,\beta,\gamma)=(1,4,1)$, which represents the Hernquist distribution, i.e. the numerical “Hernquist-type BHs”, and (ii) $(\alpha,\beta,\gamma)=(1,3,1)$, corresponding to the numerical “NFW-type BHs” where the distribution of matter follows the NFW profile. Since the total mass of the NFW profile is logarithmically divergent, we define a radial cutoff $r_{c}$, such that $M(r>r_{c})=0$ [72]. Hereafter we fix $r_{c}=5a_{0}$. Finally, to mimic the overdense cusp expected to form due to adiabatic accretion, and truncate the matter distribution at the horizon, we scale the profiles such that $\rho(r)\rightarrow\left(1-2M_{\mathrm{BH}}\right)\rho(r)$ [67, 72, 75].
@@
    $\frac{f^{\prime}(r)}{f(r)}=\frac{2m(r)/r}{r-2m(r)},\quad 2P_{t}(r)=\frac{m(r)\rho}{r-2m(r)}.$ (8)
@@ II.3 Axial perturbations
The first-order components are decomposed in terms of standard axial and polar tensor spherical harmonics [83, 84, 85] that satisfy the perturbed field equations
@@
In this work we focus on axial type perturbations, for which metric and fluid fluctuations decouple (under the assumption that there are no dissipation mechanisms in the fluid [86]). Then, axial-type gravitational fluctuations are completely governed by a single, one-dimensional, Schrodinger-like equation for a master function $\psi_{\ell m}(r)$ [87, 72]
@@
where $\ell=2,\dots,\infty$, $m=-\ell,\dots,\ell$ and $r_{*}$ is the tortoise coordinate, defined as $\mathrm{d}r_{*}/\mathrm{d}r=\left[f(r)(1-2m(r)/r)\right]^{-1/2}$. The scattering potential $V_{\ell}(r)$ reads [87, 72]:
@@
When $m(r)=M_{\mathrm{BH}}$ we recover the usual vacuum Schwarzschild background and Eq. (11) reduces to the well-known Regge-Wheeler equation [83].
@@
Equation (13) generalizes the well known results for Schwarzschild BHs [19] with a constant mass term, i.e. $m(r)=M_{\rm BH}$. Scalar and electromagnetic perturbations are described by $s=0$ and $s=\pm 1$, respectively, while for gravitational (axial) modes, $s=\pm 2$, we recover Eq. (12).
@@ III Setup for QNM calculations
This expression generalizes the boundary condition obtained both for the Schwarzschild BH and the exact Hernquist-type BH [72].
@@ III.1 Generalized matrix method for QNMs
The matrix method [88, 89, 90, 91, 92, 93, 94, 95] that is utilized and generalized here makes use of grid points in a small region of a query point to estimate its derivatives by employing a Taylor expansion. A key step of the method is to discretize the unknown eigenfunction in order to transform the differential equation and its boundary conditions into a homogeneous matrix equation. Based on the information about $N$ scattered grid points, Taylor series are carried out for the unknown eigenfunction up to $N$-th order for each grid point. The resulting homogeneous system of linear algebraic equations is solved for the eigenvalues.
Here we generalize, for the first time, the matrix method [88, 89] for generic, numerically-constructed and spherically-symmetric BH metrics, as explained in Sec. II.2. In a nutshell, when the numerical BH solutions are built, the result is a vector for each of the metric tensor components. Each vector is interpolated with high-accuracy through the standard interpolation formula, which finds a value between two points on the curve of a function. The resulting interpolated metric tensor components can then be accessed at any point of the interpolation function, thus giving us the ability to be re-discretized in order for the matrix method to work on the grid points chosen. A more analytic description for the the generalization of the matrix method is given in Appendix A.
@@
where $\tau_{0}(x),\,\lambda_{0}(x),\,\sigma_{0}(x)$ are $\omega$-dependent functions and $\phi(x)$ is a general radial function. By discretizing the interval $x\in[0,1]$, it is possible to introduce $N$ random or evenly distributed points with $x_{1}=0$ and $x_{N}=1$. Cramer’s rule allows to discretize all derivatives [88, 89] and rewrite Eq. (18) in matrix form, as $\mathcal{M}\mathcal{F}=0$, where $\mathcal{M}$ is a square matrix that depends on the eigenvalues $\omega$ of the system and $\mathcal{F}=\left(f_{1},f_{2},f_{3},\dots,f_{N}\right)^{\mathrm{T}}$ with $f_{i}=\phi(x_{i})$. The homogeneous matrix equation should satisfy $\det(\mathcal{M})=0$. The latter provides an algebraic equation to calculate the eigenvalues of any particular eigenvalue system, which can be solved numerically. The generalized boundary conditions and the generic functions $\tau_{0}(x),\lambda_{0}(x),\sigma_{0}(x)$ for asymptotically-flat, numerically-constructed BH metrics are shown in Appendix A.
Hereafter, we compute axial QNMs using $N=19$ discretization points, keeping $70$ digits of precision in all internal calculations. Convergence tests have shown that QNM frequencies converge to a given value as we increase the number of grid points, as shown in Fig. 7 of Appendix B. We also compare numerical results obtained with the generalized matrix method with those obtained through the Wentzel-Kramers-Brillouin (WKB) method. The WKB reproduces exact QNM frequencies with good accuracy in the eikonal limit, i.e. for large $\ell$. The percentage error between axial QNMs from the generalized matrix method and WKB is always, at least, less than $1\%$.
Figure 3: Top panel: Relative difference between the real part (filled dots) and the imaginary part (empty dots) of the $n=0$ fundamental axial QNMs computed for Hernquist-type, with $(\alpha,\beta,\gamma)=(1,4,1)$, and vacuum Schwarzschild BHs. We assume halo compactness between $M/a_{0}=10^{-6}$ and $M/a_{0}=10^{-2}$. Solid (and dashed) lines correspond to linear, semi-analytic fits of our data, with coefficients shown in the legend.Bottom panel: Relative percentage difference between the $n=0$ Hernquist-type BH QNM frequencies computed with the generalized matrix method and the WKB approximation, as a function of the compactness.

### v2

@@ I Introduction
For decades, black holes (BHs) have only been considered as mathematical solutions to the Einstein field equations, without any astrophysical relevance. It took many years before they were accepted as anything other than a mathematical curiosity. Even though “there should be a law of nature to prevent a star from behaving in this absurd way” [^1] Arthur Eddington, Royal Astronomical Society conference, January 1935, we now are certain that BHs not only exist but also directly affect the way we understand the cosmos in its most extreme regimes. Recent experiments detected gravitational waves (GWs) from binary BHs and neutron stars Abbott et al. (2023); Davis et al. (2021). These detections are now characterized as the pinnacle of GW astronomy Schutz (1999); Bishop (2022); Bailes et al. (2021). Combined with the foundational results of Gravity Probe A, B Vessot et al. (1980); Everitt et al. (2011), astrometry measurements at the center of our galaxy Ghez et al. (1998); Genzel et al. (2010), X-ray spectroscopy Canizares et al. (1979, 1982); Baganoff et al. (2003), infrared spectroscopy Boughn and Saulson (1983); Perlman et al. (2007) and the groundbreaking measurement of the shadows of supermassive BHs in M87* and Sgr A* with the Event Horizon Telescope Akiyama et al. (2019, 2022), the direct detection of GWs established the field of multimessenger astrophysics Mészáros et al. (2019). BHs do not only exist in nature, but also interact with other celestial objects, surrounding matter and the interstellar medium that veils the universe.
@@
Decades of simulations and observations have shown that any event involving a binary BH merger is likely to end in the same way; the GW amplitude of the merger will decay as a superposition of characteristic damped sinusoids. The damped oscillations are described by quasinormal modes (QNMs) Kokkotas and Schmidt (1999); Berti et al. (2009); Konoplya and Zhidenko (2011), i.e. a complex set of frequencies which are an intrinsic property of dissipative systems. Calculating QNMs is achieved with the use of BH perturbation theory, and has led to the establishment of the BH spectroscopy program Berti et al. (2006); Giesler et al. (2019); Baibhav et al. (2023); Destounis and Duque (2023); Berti et al. (2025). QNM frequencies are eigenmodes that are linked to the generic response of a BH under small fluctuations to the background spacetime and can help us to understand the nature of the remnant, providing precise measurements of its mass and spin, as well as allowing for tests of General Relativity (GR) Berti et al. (2006, 2009); Cardoso and Pani (2019); Franchini and Völkel (2023).
@@
The striking majority of QNM literature analyzes the spectral content of vacuum BHs, clearly an idealization. Massive BHs sit at the center of galaxies which are orders of magnitude more massive, where gaseous accretion disks Duschl (1989); Murchikova et al. (2019); Speri et al. (2023), dense dark matter spikes Daghigh and Kunstatter (2022); Chan and Lee (2024), and galactic dark matter halos form that affect the whole galaxy Wechsler and Tinker (2018). When two galaxies merge, baryonic and dark matter will interact and the supermassive BHs in their cores will meld into a new supermassive BH Struck (1999); this event is expected to emit mHz GWs and falls into the frequency band of spaceborne interferometers, like the Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017); Seoane et al. (2023); Arun et al. (2022); Karnesis et al. (2024). In this extreme case, the entire galaxy is an active participant in the dynamics and relaxation of the BH(s). The universe is also permeated with interstellar plasma, which prevents propagation of low-frequency electromagnetic waves, a screening mechanism occurs that entirely cloaks electromagnetic from GWs, which can have strong implications in BH spectroscopy Cannizzaro et al. (2024).
@@
Taking into account astrophysical environments around BHs leads to interesting effects that can affect the emission and propagation of GWs. Recently, an interesting phenomenon – spectral instability – regained a lot of interest and is directly tied to the existence of astrophysical environments around BHs. Spectral instabilities occur when QNM excessively migrate in the complex plane. This happens when environmental fluctuations, such as thin matter shells or dark matter spikes, are introduced in the vicinity of vacuum BHs. It first appeared in Refs. Nollert (1996); Nollert and Price (1999), and eventually led to the emergence of an extensive literature in relation to spectral instabilities Jaramillo et al. (2021); Cheung et al. (2022); Berti et al. (2022); Courty et al. (2023); Rosato et al. (2024); Oshita et al. (2024) and the use of non-modal tools to complement BH spectroscopy and QNMs, such as the pseudospectrum Jaramillo et al. (2022); Destounis et al. (2021); Gasperin and Jaramillo (2022); Jaramillo (2022); Boyanov et al. (2023); Destounis and Duque (2023); Destounis et al. (2024); Boyanov et al. (2024a); Cardoso et al. (2024); Areán et al. (2023); Cownden et al. (2024); Sarkar et al. (2023); Cao et al. (2024); Luo (2024). Thus, treating compact objects in non-vacuum systems will provide us with information regarding astrophysical environments and their phenomenology in GW astronomy Barausse et al. (2014).
@@
In this work, we aim in giving an extended analysis of axial QNMs for numerically-constructed BH spacetimes with different matter profiles, such as the Hernquist Hernquist (1990) and Navarro-Frenk-White (NFW) distribution Navarro et al. (1996) that are suited for different kinds of environments. Our expectation is first to provide a public code that can treat axial QNM calculations when any form of matter distribution is present, and to inspect if the redshift effect found for BHs surrounded by Hernquist-type matter Berti et al. (2025), when the compactness is varied, occurs for other matter profiles, thus extending the results regarding the detectability of astrophysical environments with axial QNMs Spieksma et al. (2024). In what follows, we use geometrized units such that $G=c=1$.
@@ II BHs in generic matter halos
Surrounding matter around perturbed astrophysical BHs can affect their GW emission and propagation. The research on non-vacuum BH QNMs Leung et al. (1997, 1999) is still in full bloom, with particular focus on galactic supermassive BHs Cardoso et al. (2022a); Konoplya (2021). Most recently, an exact general-relativistic solution of the field equations was found that describes a Schwarzschild BH embedded in a matter halo, that poses the most massive part of any galaxy. Although this approach sounds simplistic, since it does not account for luminous matter and accretion disks, to date only one exact solution of such a static spacetime has been found, where the matter halo is described by the energy density of the Hernquist profile Hernquist (1990); Cardoso et al. (2022a). In general, more spacetimes of the same nature have been constructed but only numerically or with various approximate techniques Jusufi (2023); Konoplya and Zhidenko (2022); Feng et al. (2023); Figueiredo et al. (2023); Stelea et al. (2023); Heydari-Fard et al. (2024); Speeney et al. (2024); Datta (2024); Shen et al. (2024a); Maeda et al. (2024). These fully general-relativistic configurations can, overall, provide a good description of the geometry of galaxies harboring supermassive BHs and have been proven very useful to constrain the environmental surrounding of BHs Cardoso et al. (2022a). In what follows we describe them briefly.
@@ II.1 Hernquist-type matter around static BHs
An exact, fully-relativistic solution that describes a BH immersed in a halo of matter was obtained recently Cardoso et al. (2022a). The environmental contribution in this case is accounted for through the Einstein cluster approach, which describes the matter content in terms of an anisotropic stress-energy tensor, with only tangential ($P_{t}$) and vanishing radial pressure ($P_{r}=0$), i.e. $T^{\mu}{{}_{\nu}}=\text{diag}(-\rho,0,P_{t},P_{t})$. Choosing the energy density of the fluid to follow the Hernquist distribution Hernquist (1990), i.e.,
@@
where $M_{\textrm{BH}}$ is the BH mass. The factor $1-2M_{\rm BH}/r$ in the equations above is chosen to truncate the matter at the BH horizon. It can be chosen freely, in Eq. (1) at the expense of not having a closed-form solution for the spacetime geometry. Other truncation factors, such as $1-4M_{\rm BH}/r$ will mimic matter distributions which conform better with relativistic evolutions of dark matter profiles Gondolo and Silk (1999); Sadeghian et al. (2013).
@@
The BH spacetime has an event horizon at $r=r_{h}=2M_{\textrm{BH}}$ and a curvature singularity at $r=0$. Assuming $\xi>0$ and a hierarchy of scales, i.e. $M_{\textrm{BH}}\ll M\ll a_{0}$, ensures that there are no curvature singularities outside the horizon. In the context of dark matter physics, we note that the configuration has a peak tangential pressure $P\lesssim M\rho/a_{0}$ away from the BH, thus it is Newtonian and may mimic cold dark matter well.
@@
with $U$ being the Newtonian potential of the matter distribution Duque et al. (2024). For the Hernquist distribution discussed here, $\Upsilon=2U$. We will use Eq. (6) to calculate the corrections to the light ring and its frequency, and determine the QNM frequencies in the eikonal regime.
@@ II.2 Static BHs surrounded by generic matter profiles: a numerical study
where $\rho_{0}$ is the energy density at $r=a_{0}$. The exponents $\beta$ and $\gamma$ control the slope of the distribution at large and small distances, respectively, while $\alpha$ determines the sharpness of the slope transition Merritt et al. (2006). Equation (7) motivated Refs. Figueiredo et al. (2023); Speeney et al. (2024) to develop a numerical pipeline to constructs static, and spherically symmetric BH solutions embedded in spherically symmetric, but otherwise generic, astrophysical environments.
In this work, we employ the numerical approach developed in Figueiredo et al. (2023); Speeney et al. (2024) and analyze two specific combinations of the $(\alpha,\beta,\gamma)$ parameters from Eq. (7) in order to build numerical BH solutions: (i) $(\alpha,\beta,\gamma)=(1,4,1)$, which represents the Hernquist distribution, i.e. the numerical “Hernquist-type BHs”, and (ii) $(\alpha,\beta,\gamma)=(1,3,1)$, corresponding to the numerical “NFW-type BHs” where the distribution of matter follows the NFW profile. Since the total mass of the NFW profile is logarithmically divergent, we define a radial cutoff $r_{c}$, such that $M(r>r_{c})=0$ Figueiredo et al. (2023). Hereafter we fix $r_{c}=5a_{0}$. Finally, to mimic the overdense cusp expected to form due to adiabatic accretion, and truncate the matter distribution at the horizon, we scale the profiles such that $\rho(r)\rightarrow\left(1-2M_{\mathrm{BH}}/r\right)\rho(r)$ Cardoso et al. (2022a); Figueiredo et al. (2023); Speeney et al. (2024).
@@
    $\frac{f^{\prime}(r)}{f(r)}=\frac{2m(r)/r}{r-2m(r)},\quad 2P_{t}(r)=\frac{m(r)\rho(r)}{r-2m(r)}.$ (8)
@@ II.3 Axial perturbations
The first-order components are decomposed in terms of standard axial and polar tensor spherical harmonics Regge and Wheeler (1957); Zerilli (1970); Lindblom and Detweiler (1983) that satisfy the perturbed field equations
@@
In this work we focus on axial type perturbations, for which metric and fluid fluctuations decouple (under the assumption that there are no dissipation mechanisms in the fluid Boyanov et al. (2024b)). Then, axial-type gravitational fluctuations are completely governed by a single, one-dimensional, Schrodinger-like equation for a master function $\psi_{\ell m}(r)$ Cardoso et al. (2022b); Figueiredo et al. (2023)
@@
where $\ell=2,\dots,\infty$, $m=-\ell,\dots,\ell$ and $r_{*}$ is the tortoise coordinate, defined as $\mathrm{d}r_{*}/\mathrm{d}r=\left[f(r)(1-2m(r)/r)\right]^{-1/2}$. The scattering potential $V_{\ell}(r)$ reads Cardoso et al. (2022b); Figueiredo et al. (2023):
@@
When $m(r)=M_{\mathrm{BH}}$ we recover the usual vacuum Schwarzschild background and Eq. (11) reduces to the well-known Regge-Wheeler equation Regge and Wheeler (1957). Figure 2 shows the scattering potential as a function of the radial coordinate for the Hernquist and the NFW profile, assuming $\ell=2$ and different values of the halo compactness. The inset in each plot shows the relative difference Eq. (12) and the vacuum limit.
@@
Equation (13) generalizes the well known results for Schwarzschild BHs Berti et al. (2009) with a constant mass term, i.e. $m(r)=M_{\rm BH}$. Scalar and electromagnetic perturbations are described by $s=0$ and $s=\pm 1$, respectively, while for gravitational (axial) modes, $s=\pm 2$, we recover Eq. (12).
Figure 2: Scattering potential for spin $s=\pm 2$ and $\ell=2$ for a Hernquist-type BH (upper panel) and a NFW-type BH (bottom panel) as a function of the radial coordinate $r$ normalized by the black hole mass. The absolute difference with respect to the Schwarzshild case is also shown. The compactness shown are $M/a_{0}=(10^{-6},10^{-4},\times 10^{-2})$.
@@ III Setup for QNM calculations
This expression generalizes the boundary condition obtained both for the Schwarzschild BH and the exact Hernquist-type BH Figueiredo et al. (2023).
@@ III.1 Generalized matrix method for QNMs
The matrix method Lin and Qian (2016, 2017); Lin et al. (2017); Lin and Qian (2019, 2023); Shen et al. (2023); Lin (2024); Shen et al. (2024b) that is utilized and generalized here makes use of grid points in a small region of a query point to estimate its derivatives by employing a Taylor expansion. A key step of the method is to discretize the unknown eigenfunction in order to transform the differential equation and its boundary conditions into a homogeneous matrix equation. Based on the information about $N$ scattered grid points, Taylor series are carried out for the unknown eigenfunction up to $N$-th order for each grid point. The resulting homogeneous system of linear algebraic equations is solved for the eigenvalues.
Here we generalize, for the first time, the matrix method Lin and Qian (2016, 2017) for generic, numerically-constructed and spherically-symmetric BH metrics, as explained in Sec. II.2. In a nutshell, when the numerical BH solutions are built, the result is a vector for each of the metric tensor components. Each vector is interpolated with high-accuracy through the standard interpolation formula, which finds a value between two points on the curve of a function. The resulting interpolated metric tensor components can then be accessed at any point of the interpolation function, thus giving us the ability to be re-discretized in order for the matrix method to work on the grid points chosen. A more analytic description for the the generalization of the matrix method is given in Appendix A.
@@
where $\tau_{0}(x),\,\lambda_{0}(x),\,\sigma_{0}(x)$ are $\omega$-dependent functions and $\phi(x)$ is a general radial function. By discretizing the interval $x\in[0,1]$, it is possible to introduce $N$ random or evenly distributed points with $x_{1}=0$ and $x_{N}=1$. Cramer’s rule allows to discretize all derivatives Lin and Qian (2016, 2017) and rewrite Eq. (18) in matrix form, as $\mathcal{M}\mathcal{F}=0$, where $\mathcal{M}$ is a square matrix that depends on the eigenvalues $\omega$ of the system and $\mathcal{F}=\left(f_{1},f_{2},f_{3},\dots,f_{N}\right)^{\mathrm{T}}$ with $f_{i}=\phi(x_{i})$. The homogeneous matrix equation should satisfy $\det(\mathcal{M})=0$. The latter provides an algebraic equation to calculate the eigenvalues of any particular eigenvalue system, which can be solved numerically. The generalized boundary conditions and the generic functions $\tau_{0}(x),\lambda_{0}(x),\sigma_{0}(x)$ for asymptotically-flat, numerically-constructed BH metrics are shown in Appendix A.
Figure 4: Top panel: Relative difference between the real part (filled dots) and the imaginary part (empty dots) of the $n=0$ fundamental axial QNMs computed for Hernquist-type, with $(\alpha,\beta,\gamma)=(1,4,1)$, and vacuum Schwarzschild BHs. We assume halo compactness between $M/a_{0}=10^{-6}$ and $M/a_{0}=10^{-2}$. Solid (and dashed) lines correspond to linear, semi-analytic fits of our data, with coefficients shown in the legend.Bottom panel: Relative percentage difference between the $n=0$ Hernquist-type BH QNM frequencies computed with the generalized matrix method and the 6th order WKB approximation, as a function of the compactness.
Hereafter, we compute axial QNMs using $N=19$ discretization points, keeping $70$ digits of precision in all internal calculations. Convergence tests have shown that QNM frequencies converge to a given value as we increase the number of grid points, as shown in Fig. 8 of Appendix B. We also compare numerical results obtained with the generalized matrix method with those obtained through the 6th Wentzel-Kramers-Brillouin (WKB) method. The WKB reproduces exact QNM frequencies with good accuracy in the eikonal limit, i.e. for large $\ell$. The percentage error between axial QNMs from the generalized matrix method and WKB is always, at least, less than $1\%$.